i must solve this by completing the square. 3x^2-6x=27 how do you do this
Answer by scott8148(6628)

dividing by 3 ___ x^2 - 2x = 9

adding half of the x-term coefficient, squared ___ x^2 - 2x + (-2/2)^2 = 9 + (-2/2)^2

(x - 1)^2 = 10

taking square root ___ x - 1 = ± sqrt(10)

x = 1 ± sqrt(10)